The oxidation state of bismuth in Lithium bismuthate

Grade:12

1 Answers

Souvik Mukherjee
6 years ago
Bismuthate anion has -1 ve charge, each oxygen has oxidation No. -2, now we assuming x is the Oxidation state of bismuth. Then, x + (-2) X 3 = -1, or, x – 6 = -1, or, x = 6 – 1 = +5.
 
So, the oxidation state of Bismuth in Lithium Bismuthate is +5.